A hydrate of ZnSO4 is placed in a test tube with a total mass of 54.23 g. After heating, the mass is 38.69 g. If the test tube weighs 18.83 g, what is the formula of the hydrate?
Answer #1

Mass of water = total mass - mass after heating

= 54.23 - 38.69

= 15.54 g

Moles of water = mass/molar mass = 15.54/18 = 0.863 moles

Mass of anhydrous ZnSO4 = 38.69-18.83 = 19.86 g

Moles of ZnSO4 = 19.86/161 = 0.123 moles

Molar ratio of ZnSO4 : H2O = 0.123 : 0.863

= 0.123/0.123 : 0.863/0.123

=1 : 7

thus, formula is ZnSO 4.7H2O